# Robust Bayesian Inference for Berkson and Classical Measurement Error Models

### v2

@@ 5.2 Berkson and Classical Measurement Error Studies
[…]
We first study the CASchools dataset which is available in the AER R package and provides data on various school districts in California between 1998 and 1999. Some of the features include school characteristics such as number of students, teachers and income as well as data on students’ grades in topics such as Mathematics and Reading. We employ a polynomial regression model to study the effect of school’s income on students’ test scores, as studied in Hanck et al., (2021), by considering a single outcome variable ’Score’ to be the average score between Maths and Reading. To introduce Berkson ME in the dataset we consider the case of group averages which is a common and natural way that Berkson ME arises in real world problems. This is the case when the individual measurements of a covariate are missing but we have data on the group level. People often assign to each individual the group value depending on the group they belong to and then proceed to perform inference ignoring the fact that this produces Berkson ME. To mimic this, we divide the income values range into 15 levels and we assign the within-group average income to all schools of the same income level. This creates Berkson ME in the covariates as each school in level income $i\in\{1,\dots,15\}$ has some true income $x$, but we assign to it the group average income $w_{i}$ which means that $x=w_{i}+\nu$ for some unknown error $\nu$, independent of $w_{i}$. We consider a second degree polynomial regression model, i.e.
$Y=a+bx+cx^{2}+\epsilon,\quad\epsilon\sim N(0,\sigma_{\epsilon}^{2}).$
Figure 7: Posterior model fits with pointwise 90% credible sets obtained by Robust-MEM (MMD). Scattered points indicate the observed group level average incomes (w) with associated scores (y) as well as the true unobserved individual incomes (x) with associated scores (y).
Figure 7 shows the regression fit for 200 posterior samples along with the true and observed covariates. As an oracle, we also plot a standard nonlinear Least Squares fit to the clean data, i.e. the data without ME. To assess the coverage of the suggested methods we repeat the procedure of obtaining posterior samples 100 times with different random seeds. We then compute the coverage probability as the percentage of times that the 90 % credible interval of the posterior included the true parameter values. Again, as true parameter values here we take the coefficients obtained by fitting a standard nonlinear Least Squares method to the non noisy data, i.e. the individual incomes and scores. The coverage probability for both parameters over 100 simulations was 100% at a 90% credible level whereas the average mean squared error of the residuals conditional on the true covariates over 100 runs was 0.25 with a standard deviation of 0.37.
@@
[…]
We now examine a treatment effect estimation problem in the presence of ME, in a mental health study introduced in Berry et al., (2002). The associated dataset and code for inference were later released in Harezlak et al., (2018) and code for the semiparametric Bayesian approach with penalized splines can be found in the HRW R package. The study consists of data from a six-week clinical trial of a drug versus a placebo. A physician-assessed score of the patient’s mental health is given at baseline and the score is noted again at the end of treatment. The scores are transformed and scaled and the resulting values are shown in Figure 8. Let $I_{i}$ be the indicator function for the observation $(w_{i},y_{i})$ being in the placebo group, i.e. $I_{i}=0$ if patient $i$ is in the treatment group and $I_{i}=1$ otherwise. The aim is to model the relationship between the scores at the end of treatment and at baseline, for the treatment and placebo groups, in order to estimate the average treatment effect (ATE) of the drug. The ATE here is defined as the difference between the mean response scores conditional on the true baseline scores. We denote it by:
@@
where $g^{0}(\theta^{0},x)$ and $g^{1}(\theta^{1},x)$ correspond to the mean response scores, conditional on the baseline scores, for the placebo and treatment groups respectively, indexed by parameters of interest $\theta=(\theta^{0},\theta^{1})$. Full details of the model can be found in Appendix C.3.
Following Berry et al., (2002) and Harezlak et al., (2018) we assume that the covariates $x_{i}$ are observed with ME such that for each $i=1,\dots,n$ we observe $w_{i}$ where $x_{i}\>|\>w_{i}\sim N(w_{i},0.35)$. We hence set $\operatorname{\mathbb{F}}_{W\>|\>X}=N(X,0.35)$ in the DP prior and we further set $\operatorname{\mathbb{F}}_{X}=N(0,1)$. By conjugacy we can obtain the prior centering measure $\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$ as in (11). The ATE for posterior value estimates for $\theta$ is plotted in Figure 8 with pointwise 90% credible sets along with the HRW approach denoted by BCR.
Figure 8: Left: Observed baseline scores (w) with associated end-of-treatment scores (y) for the control and treatment groups in the Mental Health study considered in Berry et al., (2002). Right: Posterior model fits for $ATE(x)=g^{1}(\theta^{1},x)-g^{0}(\theta^{0},x)$ with pointwise 90% credible sets obtained by Robust-MEM (MMD) and the BCR method of Berry et al., (2002) and Harezlak et al., (2018).
@@
To further assess our method and examine whether similar parametric assumptions might be a potential downfall, we simulate data from the obtained parameters of the BCR method and obtain posterior samples from each of the two methods. Specifically, we utilize the posterior means for each parameter obtained from the BCR method to generate new data ${(w_{i},y_{i})}_{i=1}^{n}$ according to the suggested model. In this scenario, the ground truth parameters are known, enabling us to compute the Mean Squared Error (MSE) of the residuals and the MSE of the posterior mean parameters compared to the ground truth. Table 8 presents the average MSE of the residuals across 100 simulations of the experiment, along with their associated standard deviations. Additionally, we report the MSE of the parameter estimates, where each method’s estimated parameter is taken as the posterior mean. To evaluate the uncertainty quantification of the methods, we also provide the coverage probabilities of the 90% credible posterior intervals for each parameter over 100 simulations. We observe that the suggested method outperforms BCR in terms of both residuals and parameter MSE, albeit achieving slightly lower coverage probabilities.
Remark 5.1.
Similarly to other NPL-based methods, there’s a trade-off between MSE and coverage probability due to the initialization of optimization in each bootstrap iteration, especially when the objective isn’t convex. To achieve targeted coverage probabilities, random initialization from a wide uniform distribution is effective, increasing posterior uncertainty. Conversely, for minimizing MSE, a small pre-processing step (Appendix C.3) can be used to identify an optimal point for all iterations, potentially leading to lower coverage probabilities.
@@
Table 8: Mental Health study experiment with simulated data from the suggested model with the posterior mean parameters obtained from the BCR method. New data are generated independently 100 times. Residuals MSE denotes the average MSE of the regression residuals based on the posterior mean parameters with associated standard deviation over the experiment repetitions. Parameter MSE denotes the mean squared error of the posterior mean values with respect to the true parameter value averaged over all 64 parameters.
@@ 6 Conclusion
We introduce a Bayesian NPL framework that is robust to Berkson or Classical ME and misspecified priors. Empirical findings demonstrate the robustness to ME and improved estimation accuracy of the Robust-MEM (TLS) method compared to prior art. We further instantiate our framework with the MMD-based loss enabling robust inference for nonlinear and non-Gaussian models, offering theoretical guarantees by bounding the generalization error of the Robust-MEM (MMD) method for both Berkson and Classical error models.
The proposed framework is not limited to specific parametric assumptions or availability of replicate measurements but it is flexible to incorporate such information, if it becomes available. For example, replicate measurements could be incorporated into the prior centering measures through an empirical measure. This work opens up the possibility to explore further specialized loss functions suitable for different data spaces, such as discrete, and allows the possibility to handle heteroscedastic and systematic ME scenarios via alternative nonparametric prior structures. Future work could also focus on the study of distributional properties of posterior NPL estimators, for example through extending the results of loss-likelihood bootstrap methods (Lyddon et al.,, 2019). Such results would further impact other related works based on the NPL framework.
@@ Acknowledgements
CD acknowledges support from EPSRC grant [EP/T51794X/1] as part of the Warwick CDT in Mathematics and Statistics and The Alan Turing Institute’s Enrichment Scheme. TD acknowledges support from a UKRI Turing AI acceleration Fellowship [EP/V02678X/1]. For the purpose of open access, the author has applied a Creative Commons Attribution (CC-BY) license to any Author Accepted Manuscript version arising from this submission.
@@
In Section A we provide further details on the specification, solution and optimization of the TLS problem. In Section B we provide proofs to all theoretical results presented in the main text and finally in Section C we give further experimental details on the simulations and real-world applications as well as additional experimental results.
Appendix A Further details on the TLS problem
A.1 Connection to Gaussian distributed errors
@@
which is precisely the TLS objective. If we don’t have any extra information on $\nu$, this is generally a non-identifiable problem hence TLS uses numerical optimization techniques to overcome this.
A.2 Solution of the TLS problem
Let $x\in\mathbb{R}^{n\times d_{x}}$ denote the matrix with $w_{i}\in\mathbb{R}^{d_{x}}$ as its i’th row and $Y\in\mathbb{R}^{n\times d_{y}}$ be the matrix with $y_{i}\in\mathbb{R}^{d_{y}}$ in the i’th row. Further denote by $D\in\operatorname{\mathbb{R}}^{n\times(d_{x}+d_{y})}$ the data matrix $D:=[x\quad Y]$. Let $D=u\Sigma V^{T}$ be the SVD of $D$ where $\Sigma=\text{diag}(\sigma_{1},\dots,\sigma_{d_{x}+d_{y}})$ with $\sigma_{1}\geq\dots\geq\sigma_{d_{x}+d_{y}}$ the singular values of $D$. Moreover denote by $V_{22}\in\mathbb{R}^{d_{y}\times d_{y}}$ the bottom right block of $V$. Then the TLS optimization problem in (3) has a unique solution if and only if $V_{22}$ is non-singular and $\sigma_{d_{x}}\neq\sigma_{d_{x}+1}$. The solution is given by $A=-V_{12}V_{22}^{-1}$ where $V_{12}\in\mathbb{R}^{d_{y}\times d_{x}}$ is the top right partition of $V$. A full derivation of the solution can be found at Golub and Van Loan, (1980).
Appendix B Proofs
We now provide proofs for all technical results of Section 4. We begin with all theoretical results for the Berkson case in section B.1. The proofs for the results for the Classical error case are in section B.2.
B.1 Proofs of theoretical results for the Berkson error model
B.1.1 Proof of Theorem 4.1 (Berkson)
@@ B.1 Proofs of theoretical results for the Berkson error model
Lemma B.1.
(Alquier and Gerber, (2024), Lemma S6) Let $\mathcal{S}$ be a set (equipped with a $\sigma$-algebra). Let $S_{1},\dots,S_{n}$ be independent RVs on $\mathcal{S}$ with respective distributions $Q_{1},\dots,Q_{n}$. Define $\bar{Q}=(1/n)\sum_{i=1}^{n}Q_{i}$ and $\hat{Q}=(1/n)\sum_{i=1}^{n}\delta_{S_{i}}$. Then under assumptions 4.0.1 and 4.0.2 we have:
@@
where we denote with $\mathbb{E}_{S_{1:n}\sim Q_{1:n}}$ the expectation under $S_{i}\sim Q_{i}$ for all $i=1,\dots,n$.
We now prove an important lemma bounding the expected MMD between the target marginal distribution of $X$, namely $\mathbb{P}_{X}^{0,n}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{X\>|\>w_{i}}^{0}$ where $\mathbb{P}_{X\>|\>w_{i}}^{0}$ is as defined in (18), and the empirical sample from the exact DPs $\operatorname{\mathbb{P}}_{X}$. First, denote by $\operatorname{\mathbb{F}}_{N}^{0}$ the true ME distribution, i.e. $\nu_{1:n}^{0}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}_{N}^{0}$.
Lemma B.2.
@@
$\quad\quad\quad\frac{1}{\sqrt{c+2}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})$
@@
We use the stick-breaking process representation (Sethuraman,, 1994) of a DP, meaning that for each $i\in\{1,\dots,n\}$, $\operatorname{\mathbb{P}}^{i}=\sum_{j=1}^{\infty}\xi_{j}^{i}\delta_{z_{j}^{i}}$ where $\xi_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)$ and $z_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}$. For clarity we use the separate expectations $\operatorname{\mathbb{E}}_{\xi^{i}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)}$ and $\mathbb{E}_{z^{i}_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}$ induced by $\mu_{i}$ for each $i\in\{1,\dots,n\}$. Further let $\operatorname{\mathbb{F}}^{\prime}:=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}$. Using the triangle inequality we have

### v3

@@ 6 Berkson and Classical Measurement Error Studies
We first study the CASchools dataset, available from the AER R package, which provides data on various school districts in California between 1998 and 1999. We employ a polynomial regression model to study the effect of school’s income on students’ test scores, as studied in Hanck et al., (2021), by considering a single outcome variable ‘Score’ to be the average score between maths and reading. To introduce Berkson ME in the dataset we consider the case of group averages which arises when the individual measurements of a covariate are missing but we have data on the group level. People commonly assign individuals a group value based on the group they belong to, disregarding the resulting Berkson ME. To mimic this, we divide the income values range into 15 levels and we assign the within-group average income to all schools of the same income level. This creates Berkson ME in the covariates as each school in level income $i\in\{1,\dots,15\}$ has some true income $x$, but we assign to it the group average income $w_{i}$ which means that $x=w_{i}+\nu$ for some unknown error $\nu$, independent of $w_{i}$. We consider a second degree polynomial regression model, i.e. $Y=a+bx+cx^{2}+\epsilon$ where $\epsilon\sim N(0,\sigma_{\epsilon}^{2})$. The regression fit for 200 posterior samples along with the true and observed covariates is shown in Fig. 6. As an oracle, we also plot a standard nonlinear Least Squares fit to the clean data, i.e. the data without ME. To assess the coverage of the suggested methods we repeat the procedure of obtaining posterior samples 100 times with different random seeds. We then compute the coverage probability as the percentage of times that the 90 % credible interval of the posterior included the true parameter values.
Figure 6: Posterior mean estimates with pointwise 90% credible sets obtained by Robust-MEM (MMD). Scattered points indicate the observed group level average incomes (w) and the unobserved individual incomes (x) with associated scores (y).
Again, as true parameter values here we take the coefficients obtained by fitting a standard nonlinear Least Squares method to the non noisy data, i.e. the individual incomes and scores. The coverage probability for both parameters over 100 simulations was 100% at a 90% credible level whereas the average mean squared error of the residuals conditional on the true covariates over 100 runs was 0.25 with a standard deviation of 0.37.
@@
We examine a treatment effect estimation problem in the presence of Classical ME, in a mental health study introduced in Berry et al., (2002). The associated dataset and code for inference were later released in Harezlak et al., (2018) and code for the semiparametric Bayesian approach with penalized splines can be found in the HRW R package. The study consists of data from a six-week clinical trial of a drug versus a placebo. A physician-assessed score of the patient’s mental health is given at baseline and the score is noted again at the end of treatment (Fig. 7). Let $I_{i}$ be the indicator function for the observation $(w_{i},y_{i})$ being in the placebo group, i.e. $I_{i}=0$ if patient $i$ is in the treatment group and $I_{i}=1$ otherwise. The aim is to model the relationship between the scores at the end of treatment and at baseline, for the treatment and placebo groups, in order to estimate the average treatment effect (ATE) of the drug. The ATE here is defined as the difference between the mean response scores conditional on the true baseline scores, denoted by:
@@
where $g^{0}(\theta^{0},x)$ and $g^{1}(\theta^{1},x)$ correspond to the mean response scores, conditional on the baseline scores, for the placebo and treatment groups respectively, indexed by parameters of interest $\theta=(\theta^{0},\theta^{1})$. Full details of the model can be found in Appendix E.5.
Following Berry et al., (2002) and Harezlak et al., (2018) we assume that the covariates $x_{i}$ are observed with ME such that for each $i=1,\dots,n$ we observe $w_{i}$ where $x_{i}\>|\>w_{i}\sim N(w_{i},0.35)$. We hence set $\operatorname{\mathbb{F}}_{W\>|\>X}=N(X,0.35)$ in the DP prior and we further set $\operatorname{\mathbb{F}}_{X}=N(0,1)$. By conjugacy we can obtain the prior centering measure $\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$ as in (9). The ATE for posterior value estimates for $\theta$ is plotted in Fig. 7 with pointwise 90% credible sets along with the HRW approach denoted by BCR.
Figure 7: Left: Observed baseline scores (w) with associated end-of-treatment scores (y) for the control and treatment groups in the mental health study. Right: Posterior model estimate for ATEATE with pointwise 90% credible sets obtained by Robust-MEM (MMD) and the BCR method of Berry et al., (2002) and Harezlak et al., (2018).
@@
Table 4: Simulation of mental health study using posterior mean parameters from the BCR method (100 simulations): residuals MSE (with std.), parameter MSE (MSE with std. compared to true values, averaged over 64 parameters), coverage probability (averaged over 64 parameters).
@@
To further assess our method and examine whether these assumptions might be a potential downfall, we simulate data from the obtained parameters of the BCR method and obtain posterior samples from each of the two methods. Specifically, we utilize the posterior means for each parameter obtained from the BCR method to generate new data ${(w_{i},y_{i})}_{i=1}^{n}$ according to the suggested model. In this scenario, the ground truth parameters are known, enabling us to compute the MSE of the residuals and the MSE of the posterior mean parameters compared to the ground truth. Table 4 presents the average MSE of the residuals across 100 simulations of the experiment, along with their associated standard deviations. Additionally, we report the MSE with associated standard deviations of the parameter estimates, where each method’s estimated parameter is taken as the posterior mean. We further evaluate the uncertainty quantification of the methods by providing the coverage probabilities of the 90% credible posterior intervals for each parameter over 100 simulations. We observe that the suggested method outperforms BCR in terms of both residuals and parameter MSE, albeit achieving slightly lower coverage probabilities.
Remark 6.1.
In practice there’s a trade-off between MSE and coverage due to the initialization of the optimisation step in (7). To achieve targeted coverage probabilities, random initialization from a wide uniform distribution is effective, increasing posterior uncertainty. Conversely, suppose a low MSE is a priority, then a small pre-processing step (Appendix E.5) can be used to identify an initialization point.
@@ 7 Conclusion
We introduce a Bayesian NPL framework that is robust to Berkson or Classical ME and misspecified priors. Empirical findings demonstrate the robustness to ME and improved estimation accuracy of the Robust-MEM (TLS) method compared to prior art. We further instantiate our framework with the MMD-based loss enabling robust inference for nonlinear and non-Gaussian models, offering theoretical guarantees by bounding the generalisation error of the Robust-MEM (MMD) method for both Berkson and Classical error models.
The proposed framework is not limited to specific parametric assumptions or availability of replicate measurements but it is flexible to incorporate such information, if it becomes available. For example, replicate measurements could be incorporated into the prior centering measures through an empirical measure. This work opens up the possibility to explore further specialized loss functions suitable for different data spaces, such as discrete, and allows the possibility to handle heteroscedastic and systematic ME scenarios through alternative nonparametric prior structures. Future work could also focus on the study of distributional properties of posterior NPL estimators, for example through extending the results of loss-likelihood bootstrap methods (Lyddon et al.,, 2019). Such results would further impact other related works based on the NPL framework.
@@ Acknowledgements
The authors would like to thank Mengqi Chen for spotting an error in a proof in an earlier version of the paper. CD acknowledges support from EPSRC grant [EP/T51794X/1] as part of the Warwick CDT in Mathematics and Statistics and The Alan Turing Institute’s Enrichment Scheme. TD acknowledges support from a UKRI Turing AI acceleration Fellowship [EP/V02678X/1]. For the purpose of open access, the author has applied a Creative Commons Attribution (CC-BY) license to any Author Accepted Manuscript version arising from this submission.
@@
Appendix A Outline
In Section B we give a detailed overview of related literature for inference with ME. In Section C we provide further details on the TLS problem and in Section D we provide proofs to all theoretical results presented in the main text. Finally in Section E we give further experimental details on the simulations and real-world applications as well as additional experimental results.
Appendix B Further Related Work
An overview of related work was given in Section 1.3 of the main text. There are a number of alternative approaches that have been developed to deal with ME such as Instrumental Variables methods, mainly in the Econometrics literature, Deconvolution Kernel methods, as well as methodology in the Causal Inference literature. Similarly to the approaches we discussed in the main text, these methods are often limited by strong assumptions or the requirement of additional measurements.
Instrumental Variable (IV) methods
IV methods, for example, Newhouse and McClellan, (1998); Bowden and Turkington, (1990) in the Classical setting, require access to observations from the so-called IV, which is correlated with the covariate, but not necessarily a replicate of it. The required conditional independence between the IV and the response variable, given the unobserved covariate, overcomes the bias induced by the correlation between the regression error and the covariate in the standard setting. The IV setting has also been generalised to a kernel-based nonparametric estimation of causal effect with Classical ME in the cause, in the presence of unobserved continuous confounding by Zhu et al., (2022). In the Berkson error case, Schennach, (2013) deals with nonparametric regression relying on the availability of an IV. Even in the cases where a candidate RV is available, it is hard to ensure that the necessary IV conditions hold.
Deconvolution Kernel Methods
Deconvolution kernel methods Fan and Truong, (1993); Wang and Wang, (2011) were introduced in Section 1.3 for kernel-based estimation of regression functions. One of the main challenges in these methods is the choice of the deconvolving kernel as these do not generally admit a closed form and can be sensitive to the selection of bandwidth. Lütkenöner, (2015) introduced a kernel family enabling analytical representations of deconvolving kernels for normally distributed ME. Delaigle and Hall, (2016) relaxed assumptions and proposed a deconvolution method for unknown, symmetric error distributions. Similarly, Kato and Sasaki, (2018) developed uniform confidence bands based on a multiplier-bootstrap approach, assuming unknown error distributions but requiring an independent error distribution sample. Despite desirable theore-tical properties, these methods often apply to specific function classes, such as sufficiently smooth functions. For Berkson-type error models, Delaigle et al., (2006) suggested a flexible nonparametric approach based on trigonometric functions, assuming some knowledge of the ME distribution or replicate measurements of the error or covariates.
Causal Inference
In causal inference, a biased estimate can lead to false conclusions about the causal relationship between a covariate and the response variable. This problem has been explored in the causal inference literature by reframing it as a causal effect estimation problem with unobserved confounders (see Kuroki and Pearl,, 2014; Adams et al.,, 2019; Finkelstein et al.,, 2021). Such causal effect estimation problems are met in health and epidemiology sciences where we are interested in exposure-response relations (see Brakenhoff et al.,, 2018). Exposures, often measured with error, can include both self-reported quantities, such as patient lifestyle changes, and laboratory measurements such as serum cholesterol or the levels of aluminum in the brain Campbell, (2002).
Appendix C Further details on the TLS problem
$\forall i=1,\dots,n:\quad y_{i}=\theta^{T}x_{i}+\epsilon_{i},\quad\quad x_{i}=w_{i}+\nu_{i}$
and we wish to correctly estimate $\theta\in\mathbb{R}^{d_{x}\times 1}$. It is straightforward to see that if we naively use OLS to fit $\{(w_{i},y_{i})\}_{i=1}^{n}$ with the linear model, ignoring ME, the resulting estimator for $\theta$ will be biased because each $w_{i}$ is correlated with $\nu_{i}$. This bias is not present in the TLS case. Denote by $\|\cdot\|_{F}$ the Frobenius norm, then TLS corresponds to the optimisation problem:
or equivalently:
The ordinary TLS algorithm has been extended to, among others, the regularised TLS Guo and Renaut, (2002) to induce regularisation on the set of admissible solutions and convex TLS Malioutov and Slavov, (2014) to deal with varying ME, known error structure and outliers. Some finite sample statistical properties of TLS estimators were given by Xu et al., (2012). For a more thorough review of TLS methods, we refer the reader to Van Huffel, (2004) and Markovsky and Van Huffel, (2007).
C.1 Connection to Gaussian distributed errors
@@
which is precisely the TLS objective. If we don’t have any extra information on $\nu$, this is generally a non-identifiable problem hence TLS uses numerical optimisation techniques to overcome this.
C.2 Solution of the TLS problem
The solution to the problem in (19) has been derived for real RVs using the Singular Value Decomposition (SVD) by Golub and Loan, (1979), and Golub and Van Loan, (1980) and later generalised by Van Huffel and Vandewalle, (1988) who relaxed some of the necessary numerical assumptions for the existence of a solution. Diao et al., (2019) later improved the computational time of solving the TLS problem through fast sketching methods. Here, we provide the SVD solution which we also use in the experiments.
Let $x\in\mathbb{R}^{n\times d_{x}}$ denote the matrix with $w_{i}\in\mathbb{R}^{d_{x}}$ as its i’th row and $Y\in\mathbb{R}^{n\times d_{y}}$ be the matrix with $y_{i}\in\mathbb{R}^{d_{y}}$ in the i’th row. Further denote by $D\in\operatorname{\mathbb{R}}^{n\times(d_{x}+d_{y})}$ the data matrix $D:=[x\quad Y]$. Let $D=u\Sigma V^{T}$ be the SVD of $D$ where $\Sigma=\text{diag}(\sigma_{1},\dots,\sigma_{d_{x}+d_{y}})$ with $\sigma_{1}\geq\dots\geq\sigma_{d_{x}+d_{y}}$ the singular values of $D$. Moreover denote by $V_{22}\in\mathbb{R}^{d_{y}\times d_{y}}$ the bottom right block of $V$. Then the TLS optimisation problem in (19) has a unique solution if and only if $V_{22}$ is non-singular and $\sigma_{d_{x}}\neq\sigma_{d_{x}+1}$. The solution is given by $A=-V_{12}V_{22}^{-1}$ where $V_{12}\in\mathbb{R}^{d_{y}\times d_{x}}$ is the top right partition of $V$. A full derivation of the solution can be found at Golub and Van Loan, (1980).
Appendix D Proofs
We now provide proofs for all technical results of Section 4. We begin with all theoretical results for the Berkson case in Section D.1. The proofs for the results for the Classical error case are in Section D.2.
D.1 Proofs of theoretical results for the Berkson error model
D.1.1 Proof of Theorem 4.1 (Berkson)
@@ D.1 Proofs of theoretical results for the Berkson error model
Lemma D.1.
(Alquier and Gerber, (2024), Lemma 10) Let $\mathcal{S}$ be a set (equipped with a $\sigma$-algebra). Let $S_{1},\dots,S_{n}$ be independent RVs on $\mathcal{S}$ with respective distributions $Q_{1},\dots,Q_{n}$. Define $\bar{Q}=(1/n)\sum_{i=1}^{n}Q_{i}$ and $\hat{Q}=(1/n)\sum_{i=1}^{n}\delta_{S_{i}}$. Then under assumptions 4.0.1 and 4.0.2 we have:
@@
where we denote with $\mathbb{E}_{S_{1:n}\sim Q_{1:n}}$ the expectation under $\{S_{i}\sim Q_{i}\}_{i=1}^{n}$.
We now prove an important lemma bounding the expected MMD between the target marginal distribution of $X$, namely
$\mathbb{P}_{X}^{0,n}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{P}_{X\>|\>w_{i}}^{0}$
where $\mathbb{P}_{X\>|\>w_{i}}^{0}$ is as defined in (13), and the sample from the exact DPs $\operatorname{\mathbb{P}}_{X}$. Recall that we denote by $\operatorname{\mathbb{F}}_{N}^{0}$ the true ME distribution, i.e.
$\nu_{1:n}^{0}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}_{N}^{0}$
Lemma D.2.
@@
$\quad\quad\quad\frac{1}{\sqrt{n(c+2)}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0}).$
@@
We use the stick-breaking process representation Sethuraman, (1994) of a DP, meaning that for each $i\in\{1,\dots,n\}$, $\operatorname{\mathbb{P}}^{i}=\sum_{j=1}^{\infty}\xi_{j}^{i}\delta_{z_{j}^{i}}$ where $\xi_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)$ and $z_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}$. For clarity we use the separate expectations $\operatorname{\mathbb{E}}_{\xi^{i}_{1:\infty}\sim\operatorname{\operatorname{GEM}}(c+1)}$ and $\mathbb{E}_{z^{i}_{1:\infty}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}}$ induced by $\mu_{i}$ for each $i\in\{1,\dots,n\}$. Further let $\operatorname{\mathbb{F}}^{\prime}:=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{F}}^{\prime}_{w_{i}}$. Using the triangle inequality, we have